### v1

# Crafting the Path: Robust Query Rewriting for Information Retrieval

Abstract
Query rewriting aims to generate a new query that can complement the original query to improve the information retrieval system. Recent studies on query rewriting, such as query2doc (Q2D), query2expand (Q2E) and querey2cot (Q2C), rely on the internal knowledge of Large Language Models (LLMs) to generate a relevant passage to add information to the query. Nevertheless, the efficacy of these methodologies may markedly decline in instances where the requisite knowledge is not encapsulated within the model’s intrinsic parameters. In this paper, we propose a novel structured query rewriting method called Crafting The Path tailored for retrieval systems. Crafting The Path involves a three-step process that crafts query-related information necessary for finding the passages to be searched in each step. Specifically, the Crafting The Path begins with Query Concept Comprehension, proceeds to Query Type Identification, and finally conducts Expected Answer Extraction. Experimental results show that our method outperforms previous rewriting methods, especially in less familiar domains for LLMs. We demonstrate that our method is less dependent on the internal parameter knowledge of the model and generates queries with fewer factual inaccuracies. Furthermore, we observe that Crafting The Path has less latency compared to the baselines.
Crafting the Path: Robust Query Rewriting for Information Retrieval
Ingeol Baek, Jimin Lee, Joonho Yang, Hwanhee Lee Department of Artificial Intelligence, Chung-Ang University, Seoul, Korea baekingeol98@gmail.com, {ljm1690, plm3332, hwanheelee}@cau.ac.kr
1 Introduction
In an open-domain QA system, document retrievers are utilized to retrieve the necessary information to answer the given query. Query rewriting reformulates original queries to help the retrieval system find relevant passages. Recent works on query rewriting focus on Large Language Models (LLMs) Brown et al. (2020); OpenAI (2022, 2023); Touvron et al. (2023) to generate additional information. Specifically, these studies aim to generate a relevant passage for a given query by leveraging the pre-trained knowledge of LLMs. Utilizing these new queries generated from LLMs has shown a significant increase in the performance of retrieval systems. Recently, various LLM based rewriting methods such as query2doc (Q2D) (Wang et al., 2023c), query2expand (Q2E), and querey2cot (Q2C) (Jagerman et al., 2023) have been introduced. Q2D generates a pseudo-document based on the original query, which is then used as input for the retriever. Similarly, Q2C employs a Chain-of-Thought (Wei et al., 2022) approach, and Q2E generates semantically equivalent queries. These approaches leverage the rewriting of the original query into a form similar to passages in the corpus. These techniques result in superior performance improvements compared to utilizing the base query alone.
The fundamental reason for utilizing retrieval systems in open-domain QA is to use external knowledge when QA systems do not have knowledge to generate correct answer (Lewis et al., 2020). However, relying heavily on inherent knowledge of LLMs often leads to the generation of irrelevant information and causes numerous factual errors in the reformulated queries. As in the right examples of Figure 1, Q2C asserts that coffee originated in ancient Egypt or Yemen, whereas its actual origin is Ethiopia. Meanwhile, Q2D discusses a legendary story about "goats" without providing any information related to the origin of the word "coffee". These types of misinformation and unrelated contents can lead to significant performance degradation because of the included incorrect information in the reformulated queries.
In this paper, we propose a novel query rewriting method, Crafting The Path, which is a fine-grained query reformulation technique through the structured reasoning process. Instead of simply generating the passage similar to the candidate documents, Crafting The Path focuses on identifying what information needs to be found to solve the given query. The Crafting The Path method comprises three steps. The first step, Query Concept Comprehension, provides fundamental background knowledge. Offering basic factual information reduces the likelihood of including incorrect information and helps the retrieval system clearly understand the main topic. The second step, Query Type Identification, specifies the required information to filter out irrelevant information. Finally, through Expected Answer Extraction process, the retriever model identifies the essential information it needs to find, facilitating the extraction of accurate passages. This structured, step-by-step process minimizes unnecessary inferences by the model, thereby reducing the possibility of factual errors.
Crafting The Path outperforms all of the baseline methods and exhibits a reduced degree of factual errors demonstrated by the 10% higher FActScore (Min et al., 2023) compared to baselines. Additionally, our approach demonstrates enhanced performance in the absence of prior knowledge of the model’s internal parameters. This is evidenced by experiments in closed-book QA settings, where rewriting models fail to provide correct answers, resulting in a 3.57% increase in performance. Furthermore, our method shows 7.3% less latency compared baselines.
2 Related Work
Information Retrieval
Information retrieval is the process of obtaining relevant information from the database based on given queries. The main two methods for information retrieval are sparse retrieval and dense retrieval. A prominent example of the sparse retrieval method is BM25 (Robertson and Jones, 1976; Robertson et al., 1995), which serves as a ranking function to evaluate the relevance between a given query and documents. In contrast, dense retrieval method (Xiong et al., 2021; Qu et al., 2021) involves fetching passages that exhibit high similarity to the query using the document embeddings. This approach typically utilizes pre-trained language models such as BERT (Devlin et al., 2018) for the encoder, and some methods fine-tune these encoders. (Karpukhin et al., 2020; Wang et al., 2023b). In this work, we improve the performance of both sparse and dense retrieval methods by focusing on query rewriting method.
LLM Based Query Rewriting
Query rewriting refers to the task that modifies the original query to improve the search results for the information retrieval systems. Recent studies on query rewriting mainly use the large language models to create relevant information for the given query. query2doc (Q2D) (Wang et al., 2023c) operates by generating a pseudo-document based on the original query, which is then used as input for retriever. Similarly, query2cot (Q2C) employs a Chain of Thought (CoT) (Wei et al., 2022) approach, and query2expand (Q2E) (Jagerman et al., 2023) generates semantically equivalent query. These approaches leverage the rewriting of the original query into a form similar to passages in the corpus. This technique leads to significant performance improvement compared to using the base query alone. Furthermore, Rewrite-Retrieve-Read(Ma et al., 2023) introduces a methodology that enhances rewriting performance by incorporating reinforcement learning. Another study, ITER-REGEN(Shao et al., 2023), improves query quality by feeding the query and retrieved documents into a language model for rewriting, followed by a repeated retrieval process. Rephrase and Respond (Deng et al., 2023) argues that for effective rewriting, queries should be rephrased in a manner that is easier for LLMs to understand. For the conversational serach, (Yoon et al., 2024) proposes a method that generates a variety of queries and uses the rank of retrieved passages to train the LLMs on only the optimal queries. This process is further refined using a DPO (Rafailov et al., 2023) approach to create optimal queries.
Our work has similarities with recent efforts like Q2D and Q2E, particularly in using linguistic techniques to expand queries. However, unlike previous approaches, we focus on generating only the essential information necessary for information retrieval and aim to improve performance by reducing hallucinations in the reformulated query.
3 Crafting The Path
Our proposed rewriting method, Crafting The Path, is composed of the following three steps.
3.1 Query Rewriting via Crafting The Path
3.1.1 Step 1: Query Concept Comprehension
We begin with Query Concept Comprehension step, which generates additional information that serves as the contextual background for the existing query. This step enriches the direct information about the question object within the query. For example, in Figure 1, when asking about the origin of "coffee," this step provides a detailed description of "coffee". Query Concept Comprehension step aids in identifying the information to be searched for in the next step.
3.1.2 Step 2: Query Type Identification
Based on the specific information obtained through the original query and Step 1, we proceed to Query Type Identification step. In this step, we generate the type of necessary information to retrieve relevant passages.
Specifically, we create categories for the query that help filter out irrelevant information. To retrieve information to answer the origin of coffee as in Figure 1, we can think that one must search for the historical context and etymology of coffee. Inspired by this point, Query Type Identification aims to find the type of necessary information that the ground truth passage might include as in Figure 1.
3.1.3 Step 3: Expected Answer Extraction
The final step involves extracting expected answers for the query based on the information generated from the previous step. As in Figure 1, the details regarding the origin of coffee being Ethiopia and its etymology enable the retriever model to identify the required information, facilitating the extraction of accurate passage.
| Input Prompt for Crafting The Path |
| Instruction: By following the requirements, write 3 steps related to the Query and answer in the same format as the example. |
| Requirements: |
| 1. In step1, generate the contextual background from the existing query is extracted. |
| 2. In step2, generate what information is needed to solve the question. |
| 3. In step3, generate expected answer based on query, step1, and step2. |
| 4. If you think there is no more suitable answer, end up with ’None’. |
| Query 1: what is the number one formula one car? |
| Step 1: Formula One (F1) is the highest class of international automobile racing competition held by the FIA. |
| Step 2: To know the best car, you have to look at the race records. |
| Step 3: Red Bull Racing’s RB20 is the best car. |
| (4-shot examples) … |
| Query 5: |
These three distinct steps offer a form of query rewriting that enhances the retrieval of more accurate information and minimizes the inclusion of incorrect information. We implement all of these steps in Crafting The Path with a single LLM call using the prompt in Table 1. Specifically, we provide the role of each step with the examples in the prompts. Additionally, to avoid producing inaccurate information, we instruct the model to generate “None” when it lacks certain knowledge, providing clear guidance for the retriever system’s input.
3.2 Passage Retriever
Constructing Inputs of Retriever To construct the final query , we expand the original query three times and concatenate with rewritten query in sparse retrieval as shown in Eq. 1. In the case of dense retrieval, a [SEP] token is inserted between the query and the QR to differentiate them.
Training Dense Retriever To train a dense retriever, we utilize the Binary Passage Retrieval (BPR) loss (Yamada et al., 2021) as follows to reduce the memory usage:
where denote a set where represents training instances, denotes a positive passage, and denotes a negative passage. We compute embedding using an encoder, each and represent the hash code for a query and a passage, respectively. is to identify positive passages based on ranking loss, and is the margin that is enforced between the positive and negative scores. is used to minimize the negative log-likelihood for a positive passage. Finally, we employ the BPR loss as follows:
4 Experiments
4.1 Experimental Setup
Datasets
We use the MS-MARCO passage dataset (Campos et al., 2016) for training retriever. Additionally, To demonstrate the robustness of our model on unseen data, we utilize nine retrieval datasets from BEIR (Thakur et al., 2021) for our main experiment. In our experiment, we utilize the nDCG@10 metric to evaluate the quality of the top 10 search results based on their relevance and order.
Baselines
To analyze query rewriting methods based on LLMs, we use three baselines: query2doc (Q2D), query2cot (Q2C), and query2expand (Q2E). All rewriting methods use 4-shot prompts. For Q2D, Q2E and Q2C, we reference the prompt from Wang et al. (2023c), and Jagerman et al. (2023).
In our experiments, we employ Mistral-7b (Jiang et al., 2023) and Phi-2 (Microsoft, 2023) as query rewriting models including Crafting The Path. We conduct experiments with all methods on the MS-MARCO and 9 BEIR datasets.
Implementation Details
For reliable experiments, we train five different retriever models using a different seed for each method. To evaluate the query rewriting methods using dense retrieval models, we use a total of 20 models. For each dataset, we construct 5 (fine-tuned retrieval models) 4 (rewriting methods) = 20 embedding vectors, and compute the mean and variance. In experiments with the query rewriting method using Phi-2, we obtain results using models trained with the new queries written by Mistral-7b. We employ all-mpnet-base-v2111https://huggingface.co/sentence-transformers/all-mpnet-base-v2 as our pre-trained dense retrieval model.
| scifact | trec-covid | nfcorpus | quora | scidocs | hotpotqa | dbpedia | fiqa | fever | Avg | ||
| Ours | 58.41 | 64.59 | 32.28 | 75.21 | 18.39 | 50.66 | 41.78 | 40.74 | 63.21 | 49.47 | |
| Mistral-7b | Q2D | 59.40 | 61.92 | 32.03 | 74.79 | 18.40 | 52.05 | 41.46 | 40.37 | 64.26 | 49.41 |
| Dense | Q2E | 57.06 | 55.89 | 31.98 | 71.98 | 18.29 | 48.57 | 39.13 | 40.90 | 60.85 | 47.19 |
| Q2C | 59.37 | 63.14 | 32.15 | 74.33 | 18.31 | 52.27 | 41.75 | 40.60 | 62.60 | 49.39 | |
| Ours | 58.40 | 63.24 | 32.60 | 75.44 | 17.99 | 49.84 | 40.32 | 40.31 | 62.12 | 48.92 | |
| Phi-2 | Q2D | 58.23 | 59.50 | 31.52 | 75.55 | 18.05 | 49.28 | 39.68 | 39.81 | 62.73 | 48.26 |
| Dense | Q2E | 57.07 | 56.58 | 31.47 | 72.46 | 18.26 | 46.94 | 38.44 | 40.61 | 59.84 | 46.85 |
| Q2C | 57.84 | 62.23 | 32.20 | 73.56 | 18.00 | 49.39 | 40.20 | 39.61 | 60.95 | 48.22 | |
| Ours | 70.78 | 74.79 | 35.51 | 75.82 | 16.10 | 57.41 | 46.68 | 29.12 | 63.66 | 52.21 | |
| Mistral-7b | Q2D | 71.14 | 67.73 | 35.01 | 70.37 | 15.68 | 58.82 | 41.36 | 28.67 | 67.72 | 50.72 |
| Sparse (BM25) | Q2E | 68.61 | 66.29 | 35.14 | 76.66 | 16.13 | 54.58 | 42.69 | 28.15 | 52.98 | 49.03 |
| Q2C | 71.63 | 74.45 | 35.29 | 74.86 | 16.10 | 58.71 | 43.16 | 30.91 | 64.36 | 52.16 | |
| Ours | 68.48 | 69.71 | 34.38 | 74.47 | 15.47 | 55.20 | 42.93 | 28.91 | 59.20 | 49.86 | |
| Phi-2 | Q2D | 67.15 | 59.42 | 32.83 | 71.45 | 14.95 | 52.61 | 37.50 | 25.10 | 57.67 | 46.52 |
| Sparse (BM25) | Q2E | 68.71 | 66.94 | 33.51 | 77.35 | 16.21 | 53.35 | 39.18 | 26.72 | 52.53 | 48.28 |
| Q2C | 69.14 | 69.57 | 35.21 | 75.33 | 15.70 | 55.39 | 39.64 | 28.37 | 59.57 | 49.77 | |
4.2 Results
Main Results
As shown in Table 2, Crafting The Path outperforms all approaches, including Q2D, Q2C, and Q2E, for the average score across 9 BEIR datasets. Our approach consistently demonstrates superior performance compared to existing methods. Crafting The Path provides a robust application across various domains and rewriting model sizes in both sparse and dense retrievers. Moreover, we observe significant performance improvements on the trec-covid (Wang et al., 2020; Voorhees et al., 2020) and nfcorpus (Boteva et al., 2016) datasets for both retriever types compared to previous methods. Especially, trec-covid dataset requires searching for the latest information on queries about COVID-19. Crafting The Path proves to be more effective in finding such recent information. However, previous methods outperform our approach on datasets like FEVER (Thorne et al., 2018). The FEVER dataset is based on Wikipedia, which is frequently used for LLMs pre-training data. Since the training data for the Mistral-7b and Phi-2 models are not disclosed, and verifying the presence of internal parameter knowledge directly remains a challenge, it is difficult to confirm the internal knowledge of the models (Wang et al., 2023a). However, we conduct measurements of the impact of internal model parameters through experiments in Table 5. As results in Table 2, we observe the significant performance improvements in these BEIR datasets from Crafting The Path rewriting methods through the use of a structured rewriting method.
Ablation Study
We conduct an ablation study on Crafting The Path. We measure the performance on the MS-MARCO dataset by excluding each step as shown in Table 3. We experiment with omitting Step 3 and also conduct experiments excluding both Steps 2 and 3. Performance declines with the removal of each step, demonstrating that each step is essential for performance improvement.
| MS-MARCO Passage dev | |||
| nDCG | MRR | Recall@1K | |
| Crafting The Path | 45.42 | 33.11 | 97.05 |
| w/o step3 | 44.99 | 32.91 | 96.67 |
| w/o step2, 3 | 44.59 | 32.42 | 96.17 |
| BEIR | FActScore | ||
| Crafting the Path | Q2D | Q2C | |
| Mistral-7b | 0.718 | 0.506 | 0.711 |
| Phi-2 | 0.765 | 0.460 | 0.675 |
4.3 Analysis
Evaluating Factuality of Queries
To determine the impact of factual errors occurring during query rewriting on retrieval performance, we use FActScore to measure the accuracy of Mistral-7b and Phi-2 across three rewriting methods. Unlike Min et al. (2023), which uses atomic facts, we simply divide the content by sentence. For each separated sentence, we use a gold label passage as evidence to output as True or False. If the factuality evaluation results for three sentences of a rewritten query are True, True, False, we assign a score of 2/3. We calculate the average for the measured queries. In Table 4, Crafting The Path exhibits the smallest factual error, demonstrating the best performance with both the Mistral-7b and Phi-2 models.
| hotpotQA | Correct Answer | Incorrect Answer | ||||
| Mistral-7b | Ours | Q2D | Q2C | Ours | Q2D | Q2C |
| nDCG@10 | 77.80 | 76.00 | 79.27 | 60.05 | 55.57 | 58.77 |
| MRR | 90.04 | 87.73 | 90.90 | 78.76 | 71.71 | 75.35 |
| Phi-2 | Ours | Q2D | Q2C | Ours | Q2D | Q2C |
| nDCG@10 | 74.72 | 71.03 | 75.43 | 59.42 | 46.93 | 57.64 |
| MRR | 87.95 | 86.17 | 87.75 | 79.16 | 64.63 | 75.59 |
Measuring the Reliance on Internal Knowledge
To evaluate the reliance of the internal model parameter knowledge in query reformulation of each LLM, we divide the dataset into problems where each LLM can generate the correct answer and those where it cannot in a closed-book setting. Based on this division, we apply three rewriting methods and measure scores as shown in Table 5. Both the Mistral-7b and Phi-2 models demonstrate superior performance in the Incorrect Answer cases when using our rewriting method, compared to previous rewriting approaches. This demonstrates that our method achieves more effective information retrieval when the model needs to search unknown information, aligning with the original purpose of using RAG (Lewis et al., 2020). In this experiment, we use the Contriever (Izacard et al., 2022) model for the dense retriever.
Answer Modification
In Table 6, rewriting the HotpotQA dev dataset (Yang et al., 2018), comprising 5,447 entries using Mistral-7b, results in new queries that include 1,746, 1,623, and 1,752 answers for Crafting The Path, Q2D, and Q2C, respectively. To evaluate the impact of our model’s reliance on internal knowledge for generating answers, we conducted three experiments. First, in the New Query setting, we directly utilize the new queries generated by the LLM. Second, in the Replace Answer to [MASK] setting, if the new query contained an answer, we mask the answer portion. Finally, in the Delete New Queries with Answer experiment, we exclude any new queries from the evaluation if at least one of the three rewriting methods generated a new query containing an answer. The evaluation dataset reduces to 3,212 in Delete New Queries with Answer experiment. Our approach achieves the best performance across all three experiments. Notably, the difference in nDCG@10 scores between our method and Q2C was 0.45 in the first experiment and 1.75 in the third experiment. This score gap suggests that our approach is relatively less affected by the presence or absence of answers in the queries compared to existing methods.
| hotpotqa | ||
| Mistral-7b | nDCG@10 | MRR |
| New Query | ||
| Crafting The Path | 65.49 | 82.21 |
| query2doc (Q2D) | 61.83 | 76.62 |
| query2cot (Q2C) | 65.04 | 80.11 |
| Replace Answer to [MASK] | ||
| Crafting The Path | 64.09 | 81.45 |
| query2doc (Q2D) | 61.08 | 76.36 |
| query2cot (Q2C) | 63.92 | 79.59 |
| Delete New Queries with Answer | ||
| Crafting The Path | 58.29 | 77.46 |
| query2doc (Q2D) | 53.39 | 69.91 |
| query2cot (Q2C) | 56.54 | 73.68 |
| LLM call | Index search | |
| Crafting The Path | 5648.9ms | 261.44ms |
| query2doc (Q2D) | 8167.5ms | 335.65ms |
| query2cot (Q2C) | 6094.2ms | 281.98ms |
Latency
Since all three methods use the same dense retriever architecture, the search time is equal. Hence, we measure the time of LLM call latency to compare the speed of each method. We retrieve results from the MS-MARCO passage dev dataset for 1,000 entries and average the outcomes over 100 repetitions. We measure the latency incurred when the model generates a new query and BM25 searches for relevant passages. Our method generates less redundant information, resulting in lower latency compared to Q2D. Also, our method shows comparable latency compared to Q2C due to the similar length of generated text.
5 Conclusion
We present a Crafting the Path, an approach that involves a structured three-step process, focusing not merely on generating additional information for the query but primarily on generating what information to find from the query. Our method less relies on internal model parameters and demonstrates robust performance across various models. Additionally, the method generates fewer factual errors and delivers improved out-of-domain performance with lower latency than previous methods.
Acknowledgement
This research was supported by Institute for Information & Communications Technology Planning & Evaluation (IITP) through the Korea government (MSIT) under Grant No. 2021-0-01341 (Artificial Intelligence Graduate School Program (Chung-Ang University)).
Appendix A Appendix
A.1 Hyperparameters
| Name | Value |
| Learning rate | 2e-5 |
| PLM | all-mpnet-base-v2 |
| Batch | 128 |
| Epoch | 3 |
| Learning rate decay | linear |
| Warmup steps | 1000 |
| Binary loss margin | 2.0 |
| Similarity function | dot score |
| Query length | 128 |
| Passage length | 128 |
A.2 Prompts
| prompts |
| prompts |
| prompts |
| prompts |